OpenStudy (anonymous):

Solve. z + 17 = 9 A. 9 B. –9 C. 8 D. –8

OpenStudy (anonymous):

i think c

OpenStudy (anonymous):

isolate x. z+17=9 -17 -17 z=-8

OpenStudy (anonymous):

close

OpenStudy (anonymous):

thanks